# Image Classification with Rotation-Invariant Variational Quantum Circuits

### v1

@@ I Introduction
Together with the recent surge of quantum technologies, both in theoretical frameworks and practical applications, quantum computing has been placed in the spotlight. Many quantum algorithms have shown significant speed-ups with respect to classical algorithms, and a wide variety of applications are being explored, from simulation of many-body systems to combinatorial optimization, cryptanalysis, database search, machine learning, finance, and beyond [1, 2, 3, 4, 5, 6, 7, 8]. However, many of these algorithms require hardware resources that stand far from the capabilities of the currently available noisy intermediate-scale quantum (NISQ) devices. For this reason, and in parallel to foundational algorithmic research focused on a quantum advantage in fault-tolerant quantum devices, a significant effort is put into the search for mid-term applicability and utility of algorithms in the NISQ era. Although its computational advantage is yet to be proven, Quantum Machine Learning (QML) seems one of the most promising candidates for an early application of quantum devices [9, 10, 11], since in many algorithms the goal is to approximate a distribution and the result often does not rely on an exact value of the computation. This fact makes QML algorithms resilient to noise, and some signs hint that the noise could even improve the optimization in specific cases [12, 13].
@@
QML encompasses a broad meaning, but the scope of this work will focus on the employment of quantum circuits in learning problems that involve classical data. Variational quantum algorithms [14, 15, 16, 17] hold a renowned place in this paradigm, where quantum circuits are used to create expressive high-dimensional feature maps that have been proved to be asymptotically universal approximators [18, 19]. However, these variational algorithms suffer trainability issues called Barren Plateaus (BP), where the variance of the loss function tends to vanish exponentially with the size of the quantum circuit [20, 21]. Extensive work has been done characterizing this phenomenon, identifying its sources [22, 23, 24, 25, 26] and looking for methods to mitigate or avoid it [27, 28, 29].
@@
Geometric Quantum Machine Learning (GQML) [30, 31, 32, 33], has shown a great potential to alleviate Barren Plateaus and bring significant improvements in trainability and generalization of Variational Quantum Classifiers [34]. Classical Geometric Deep Learning [35] founded an algebraic framework to analyze learning problems and proposed a geometric approach to add inductive bias [36] to learning models, i.e., a blueprint to design algorithms based on the prior knowledge about the geometric properties of the data set. As its quantum counterpart, GQML has established a potential research line where many authors implemented the geometric approach to a variety of practical problems, showing promising results. For example, Ref.[37] studied the behavior of equivariant quantum neural networks (EQNN) under the presence of noise and suggested a strategy to enhance the protection of symmetry; Ref.[38] introduced an EQNN for a classification task with $\mathbb{Z}_{2}$ permutation symmetry; in Ref.[39] Quantum Fourier Transform was employed to build a rotation equivariant QNN for scanning tunneling microscope images; Ref.[40] built a $SU(2)$- equivariant quantum circuit to learn on spin networks; Ref.[41] proposed a data-encoding strategy for Point Clouds, which is invariant under point permutations; and Ref.[42] designed a permutation-equivariant quantum circuit to solve the Traveling Salesman Problem and train it using Reinforcement Learning. Related to this work, Ref.[43] presented an equivariant quantum classifier, in this case, for images with reflection label-symmetry. Additionally, Ref.[44] covered the case of images with $p4m$ label-symmetry.
@@
In this context, the aim of this paper is to introduce an architecture for a variational quantum circuit for image classification, equivariant to $C_{4}$ image-rotation label-symmetry. In difference to the work of Ref.[43] and Ref.[44], which both of them used Ampitude Encoding, this work employs a different encoding strategy, namely Angle Encoding as discussed in III.2. This encoding allows the application of re-uploading layers, boosting the approximation ability of the model. Moreover, reflection symmetry is not included in our scope, making the proposed model distinguish chiral images. This makes the model fit, for example, for medical vision problems like brain tumor image detection or fingerprint classification, where it is vital to discern between the left and right sides of the body. Concretely, the contributions of this work are as follows:
@@ II.2 Variational Quantum Classifier
The outputs of the quantum circuit are then evaluated in the loss function, where a classical optimizer will determine the update of variational parameters, making use of techniques such as parameter shift [46] to compute the gradient of the loss function respect to the optimizable parameters.
@@ II.3 Geometric Quantum Machine Learning
It is said that a data set contains a symmetry if a certain property of the data set is preserved under the action of an operation, namely a symmetry operation. The set of all such operations $g$ along with a composition operation between them, constitute a group $G$ [47]. It is important to keep in mind that this symmetry group $G$ is defined in general and abstract terms, and it can act over different mathematical structures through its different representations. To illustrate this, it can be thought of two representations of the same data-point linked by a data encoding unitary $U(x)$. This unitary maps a data-point from its original space $\mathcal{X}$ to the Hilbert space $\mathcal{H}$ of our quantum circuit. The symmetry operation $g\in G$ then will have two representations acting on each one of the spaces the data-point is represented:
@@
where $V_{g}$ is the representation of $g\in G$ acting on the original data-space and $U_{g}$ its representation acting on the Hilbert’s space. Quantum gates are represented as unitary square matrices from the group $U(2^{n})$, where $n$ stands for the number of qubits they act on. In consequence, it is possible to represent $U_{g}$ as a $2^{n}\times 2^{n}$ unitary matrix, but it is remarkable how its ultimate representation will be determined by the choice of the data encoding method $U(x)$.
The purpose of this framework is to provide some geometric tools to design label-invariant VQCs respect to the action of $G$:
@@
being $f$ the output response of the model. To achieve this goal, the first step is to ensure that the chosen encoding unitary is equivariant respect to the symmetry group [32]:
@@
which, applied to a unitary $A$, returns another unitary that commutes with all symmetry operations in the symmetry group $[\mathcal{T}(A),U_{g}]=0\;\;\forall g\in G$. Accordingly, if the variational unitary $W$ is built out from symmterized gates, it will satisfy $[W(\theta),U_{g}]=0$, and all re-uploading layers $U_{l}(x,\theta_{l})=U(x)W_{l}(\theta_{l})$ will be equivariant respect the symmetry group.
Finally, if an invariant initial-state $|0\rangle=U_{g}|0\rangle$ and an invariant observable $O=U_{g}OU_{g}^{\dagger}$ are chosen, both for $\forall g\in G$, a label-invariant VQC will be obtained:
$f(V_{g}[x],\theta)$ $=$ $\langle 0|U^{\dagger}(V[x])W^{\dagger}(\theta)OW(\theta)U(V[x])|0\rangle$ (8)
$=$ $\langle 0|U_{g}U^{\dagger}(x)U_{g}^{\dagger}W^{\dagger}(\theta)OW(\theta)U_{g}U(x)U_{g}^{\dagger}|0\rangle$
@@ III.1 Synthetic toy dataset
The employed data set will be formed by square images of $n\times n$ pixels, where each pixel will take values laying between 0 and 255 in grayscale. Hence, each image can be represented by a matrix $x\in GL_{n}(\mathbb{R})$. For a first stage, images with resolution $n=4$ of T and L tetrominoes have been generated (see fig.2). Each tetromino can be placed in 6 different positions and 4 orientations inside an image of $4\times 4$ pixels, but it is possible to augment the size of the data set arbitrarily by adding a layer of noise to each image.
@@
A crucial step to build a label-invariant quantum model is to find the representation of this operation acting on the Hilbert space of the quantum circuit. To do so, first, the chosen data-encoding strategy will be described.
@@ III.2 Angle encoding
Frequently named as qubit encoding, angle encoding [11, 48] is one of the most simple data-encoding methods. It scales linearly in the number of required qubits and can be implemented in constant time, making it one of the most fit candidates to apply on NISQ devices. It consists on encoding the features $x_{k}$ of the data point $x=(x_{0},\dotsi,x_{N})$ in the amplitudes of the state-vector of each qubit:
$|x\rangle=\bigotimes_{k=0}^{N}\cos(x_{k})|0\rangle+\sin(x_{k})|1\rangle.$ (10)
This encoding can be performed by applying $R_{X}(x_{k})$ gates over each qubit (see fig.3), where the pixel $x_{i,j}$ is mapped to $x_{k}$ vector value with $k(i,j)=in+j$ index transformation.
@@
Due to the periodic nature of $R_{X}$ rotation gates, the classical features ought to be scaled to fit in $x_{k}\in[-\pi,\pi]$. This method requires one qubit per feature in the data-point, which can be reduced by a factor of two encoding an additional feature in the complex phase of each qubit, following the so-called dense angle encoding strategy. This improvement in the number of required qubits $n_{q}$, which is still linear with respect to the number of features $\mathcal{O}(N)$, is notably surpassed by the exponential reduction achieved with amplitude encoding [49] $\mathcal{O}(\log_{2}(N))$. Nevertheless, this improvement in the number of qubits comes along with an exponential circuit depth and the impossibility of performing data re-uploading. For this reason, the angle encoding method will be employed in this work and a way to handle the issue of the number of qubits will be discussed in subsequent stages.
@@ III.3 Equivariant variational circuit
Figure 4: Rotation operation decomposed into a (3(a)) transposition and (3(b)) column permutation using SWAP operations. Example for an image of n=4.
@@
$\mathcal{T}[X_{k}]$ $=$ $\frac{1}{4}(X_{0}+U_{g}X_{k}U_{g}^{\dagger}+U_{g}^{2}X_{k}{U_{g}^{\dagger}}^{2}+U_{g}^{3}X_{k}{U_{g}^{\dagger}}^{3})$ (19)
@@
Let orbit $\varphi$ denote the set of all different indices obtained under the action of all symmetry operations for a given pair of indices $(i,j)$:
@@
Consequently, the same symmetrization is obtained for rotation gates with the same rotation angle applied on qubits connected by the same orbit. This symmetrization is equivalent for the Y and Z rotation axes. The same logic can be followed for two-qubit gates, leading to this result: the same two-qubit gate must be applied over the orbits of both qubits, which must be in different orbits, to ensure the permutation condition of these gates with respect to the representation of the symmetry group. To illustrate this, as a reference to condensed matter physics, we will call unit cell to a set of indices that don’t share orbits; this is, a set of points that are not related by a symmetry operation (see fig. 5).
@@
Figure 5: A schematic image of different orbits (4(a)) and different unit cells (4(b), 4(c)) represented by colors for an image of $n=4$. 4(b) and 4(c) show the two-qubit gate architecture chosen to build an equivariant variational unitary for an image of $n=4$. It is built out of intra-cell and inter-cell $CNOT$ gates.
Apart from the constraints concluded from the symmetrization of two-qubit gates, one has a certain freedom to choose their connectivity, i.e., on which qubit it can apply them. Indeed, the choice of the unit cell shown in 4(b) and 4(c) is not unique. A circular connectivity scheme will be selected,which entangles the qubits inside every unit cell. However, if different unit cells are not connected between them, it will produce four independent quantum circuits that would process just some phew pixels of the original image. To avoid this scenario, some CNOTs can be added between unit cells and deal with a more extensive set of correlated qubits.
@@ III.4 Simulation results
An equivariant quantum circuit has been trained to learn to classify tetromino images described in III.1. This circuit employs parameterized quantum gates symmetrized using the Twirling Formula, following the workflow detailed in III.3. In consequence, a label-invariant model was obtained with respect to the action of 4-fold rotations. To benchmark the power of the model and to witness whether the geometric approach does provide any advantage in comparison to conventional variational circuits, its performance was compared to a widely used architecture called Basic Entangler Layers, based on [50]. To make the comparison the fairest possible and since the latter architecture makes use of more optimizable parameters per layer, which can induce an improved approximation capacity, general rotation gates were used in the Equivariant model, which can be decomposed into $3$ optimizable rotation gates, up to a negligible global phase [51]:
@@
It is true that these general rotation gates are more expressive than one-axis rotation gates employed in the Basic Entangler model, and this could be a possible source of a difference in performance. For this reason, a third model type was compared, which we will call Non-equivariant model. Its quantum circuit utilizes the same architecture as the Equivariant model’s but with randomly generated orbits to break the symmetrization of gates and, consequently, the equivariant condition. This way, we ensure we capture if any improvement in the performance is due to the choice of gates or if it is, indeed, a consequence of the geometric approach. See tab. 1 to compare the number of optimizable parameters in each model.
Model Parameters
Equivariant $3n_{\varphi}n_{l}=12n_{l}$
Non-equivariant $3n_{\varphi}n_{l}=12n_{l}$
Basic Entangler $n_{q}n_{l}=16n_{l}$
@@
Table 1: The number of optimizable parameters in the quantum circuits. $n_{\varphi}$ corresponds to the number of different orbits in an image, $n_{l}$ is the number of data re-uploading layers, and $n_{q}$ corresponds to the number of qubits in the circuit. For an image of $n=4$, $n_{\varphi}=4$ and $n_{q}=16$.
@@
Figure 6: Loss landscape of three model types along the training process where $10$ models were trained for each type. The shaded areas are spread between the maximum and minimum costs of each model type, and the dashed lines represent the mean cost of each model type. All models apply $n_{l}=5$ re-uploading layers.
@@
All three models were trained in a noiseless environment using the default simulator provided by Pennylane [52], and the parameter update is performed using Adam Optimizer [53]. In fig. 6, it can observed that the Equivariant model reaches the lowest values in the loss landscape for the same number of iterations. Even if the Basic Entangler performs a steeper descent in early steps, it reaches a plateau where the cost nearly stops decreasing. The other two models, however, keep reducing the cost until the last optimization step, although their descent tends to flatten in the last iterations.
@@
Figure 7: Violin plots of F1 score metrics of trained models. Ten different models of each type were trained with the same hyperparameters ($n_{l}=5$), and the performance was evaluated over the training and test set.
@@
Table 2: Performance metrics for each model type and per different number of data re-uploading layers. Each value in the table represents the mean performance metric of 5 different models, evaluated over the training set (left) and the test set (right).

### v2

@@ I Introduction
Together with the recent surge of quantum technologies, both in theoretical frameworks and practical applications, quantum computing has been placed in the spotlight. Many quantum algorithms have shown significant speed-ups with respect to classical algorithms and a wide variety of applications are being explored, from simulation of many-body systems to combinatorial optimization, cryptanalysis, database search, machine learning, finance, and beyond [1, 2, 3, 4, 5, 6, 7, 8]. When talking about speed-ups or quantum advantage, it is referred to proven algorithmic improvement of the computational complexity needed to solve a certain problem, compared to the best known classical algorithm [9, 10, 11]. However, many of these algorithms require hardware resources that stand far from the capabilities of the currently available noisy intermediate-scale quantum (NISQ) devices. For this reason, and in parallel to foundational algorithmic research focused on a quantum advantage in fault-tolerant quantum devices, a significant effort is put into the search for mid-term applicability and utility of algorithms in the NISQ era. Although its computational advantage is yet to be proven, Quantum Machine Learning (QML) seems one of the most promising candidates for an early application of quantum devices [12, 13, 14], since in many algorithms the goal is to approximate a distribution and the result often does not rely on an exact value of the computation. This fact makes QML algorithms resilient to noise, and some signs hint that the noise could even improve the optimization in specific cases [15, 16].
@@
QML encompasses a broad meaning, but the scope of this work will focus on the employment of quantum circuits in learning problems that involve classical data. Variational quantum algorithms [17, 18, 19, 20] hold a renowned place in this paradigm, where quantum circuits are used to create expressive high-dimensional feature maps that have been proved to be asymptotically universal approximators [21, 22]. However, these variational algorithms suffer trainability issues called Barren Plateaus (BP), where the variance of the loss function tends to vanish exponentially with the size of the quantum circuit [23, 24, 25]. Extensive work has been done characterizing this phenomenon, identifying its sources [26, 27, 28, 29, 30, 31, 32, 33] and looking for methods to mitigate or avoid it [34, 35, 36].
@@
Geometric Quantum Machine Learning (GQML) [37, 38, 39, 40], has shown a great potential to alleviate Barren Plateaus and bring significant improvements in trainability and generalization of Variational Quantum Classifiers [41]. Classical Geometric Deep Learning [42] founded an algebraic framework to analyze learning problems and proposed a geometric approach to add inductive bias [43] to learning models, i.e., a blueprint to design algorithms based on the prior knowledge about the geometric properties of the data set. As its quantum counterpart, GQML has established a potential research line where many authors implemented the geometric approach to a variety of practical problems, showing promising results. For example, Tüysüz, Cenk, et al. studied the behavior of equivariant quantum neural networks (EQNN) under the presence of noise and suggested a strategy to enhance the protection of symmetry [44]; Zheng, Han, et al. introduced an EQNN for a classification task with $\mathbb{Z}_{2}$ permutation symmetry [45]; East, R. D., et al. built a $SU(2)$- equivariant quantum circuit to learn on spin networks [46]; Heredge, Jamie, et al. proposed a data-encoding strategy for Point Clouds, which is invariant under point permutations [47]; and Skolik, Andrea, et al. designed a permutation-equivariant quantum circuit to solve the Traveling Salesman Problem and train it using Reinforcement Learning [48].
All this work contributes significantly to the state-of-the-art of GQML, but only a few articles can be found that apply the geometric approach to the problem of image classification, which is one of the main tasks in supervised learning. Among them, West, M. T., eta al. presented an equivariant quantum classifier for images with reflection label-symmetry [49] and a rotation equivariant Quantum Neural Network for scanning tunneling microscope images, employing Quantum Fourier transform [50]. Additionally, Chang, S. Y., et al. proposed an equivariant Quantum Convolutional Neural Network for images with $p4m$ label-symmetry [51]. Nevertheless, all these works employ the same data encoding method, namely amplitude encoding, which offers an exponential qubit reduction, but at the same time, a circuit depth overhead that scales exponentially with the number of qubits [14, 52, 53, 54]. On the other hand, a gap is identified in the literature, covering the application of GQML to image classification, employing angle encoding, which allows an efficient application of re-uploading layers, boosting the approximation ability of the model arbitrarily with the number of applied layers. This encoding can be implemented in linear depth (see Sec.III.2), and thus, can be executed in NISQ devices.
@@
In this context, this paper aims to introduce an architecture for a variational quantum circuit for image classification, equivariant to $C_{4}$ image-rotation label-symmetry, and using angle encoding. The simple choice of the data encoding scheme determines completely the gate symmetrization of the quantum circuit, which makes the proposed architecture differ from the aforementioned related work. In this case, the $C_{4}$ symmetry group is considered, but more symmetry operations, such as image reflection symmetry, can be taken into account following the same methodology of gate symmetrization. However, reflection symmetry is not included in the scope, making the proposed model distinguish chiral images. This makes the model fit, for example, for medical vision problems like brain tumor image detection or fingerprint classification, where it is vital to discern between the left and right sides of the body. Additionally, since one of the bottleneck of angle encoding is the required number of qubits for the codification, equivariant classical convolution layers are proposed to reduce the dimensionality of the data, while preserving the symmetry properties of the whole model. This enables the quantum model to handle large resolution images.
Concretely, the contributions of this work are as follows:
@@ II.2 Variational Quantum Classifier
The outputs of the quantum circuit are then evaluated in the loss function, where a classical optimizer will determine the update of variational parameters, making use of techniques such as parameter shift [56] to compute the gradient of the loss function with respect to the optimizable parameters.
@@ II.3 Geometric Quantum Machine Learning
It is said that a data set contains a symmetry if a certain property of the data set is preserved under the action of an operation, namely a symmetry operation. The set of all such operations $g$ along with a composition operation between them, constitute a group $G$ [57, 58, 59]. It is important to keep in mind that this symmetry group $G$ is defined in general and abstract terms, and it can act over different mathematical structures through its different representations [38]. To illustrate this, it can be thought of two representations of the same data-point linked by a data encoding unitary $U(x)$. This unitary maps a data-point from its original space $\mathcal{X}$ to the Hilbert space $\mathcal{H}$ of our quantum circuit. The symmetry operation $g\in G$ then will have two representations acting on each one of the spaces the data-point is represented:
@@
where $V_{g}$ is the representation of $g\in G$ acting on the original data-space and $U_{g}$ its representation acting on the Hilbert’s space. Quantum gates are represented as unitary square matrices from the group $U(2^{n})$, where $n_{q}$ stands for the number of qubits they act on. In consequence, it is possible to represent $U_{g}$ as a $2^{n_{q}}\times 2^{n_{q}}$ unitary matrix, but it is remarkable how its ultimate representation will be determined by the choice of the data encoding method $U(x)$.
The purpose of this framework is to provide some geometric tools to design label-invariant VQCs with respect to the action of $G$:
@@
being $f$ the output response of the model. To achieve this goal, the first step is to ensure that the chosen encoding unitary is equivariant with respect to the symmetry group [39]:
@@
which, applied to a unitary $A$, returns another unitary that commutes with all symmetry operations in the symmetry group $[\mathcal{T}(A),U_{g}]=0\;\;\forall g\in G$. Accordingly, if the variational unitary $W$ is built out from symmterized gates, it will satisfy $[W(\theta),U_{g}]=0$, and all re-uploading layers $U_{l}(x,\theta_{l})=U(x)W_{l}(\theta_{l})$ will be equivariant with respect to the symmetry group.
Finally, if an invariant initial-state $|\psi_{0}\rangle=U_{g}|\psi_{0}\rangle$ and an invariant observable $O=U_{g}OU_{g}^{\dagger}$ are chosen, both for $\forall g\in G$, a label-invariant VQC will be obtained:
$f(V_{g}[x],\theta)$ $=$ $\langle\psi_{0}|U^{\dagger}(V[x])W^{\dagger}(\theta)OW(\theta)U(V[x])|\psi_{0}\rangle$ (8)
$=$ $\langle\psi_{0}|U_{g}U^{\dagger}(x)U_{g}^{\dagger}W^{\dagger}(\theta)OW(\theta)U_{g}U(x)U_{g}^{\dagger}|\psi_{0}\rangle$
@@ III.1 Synthetic toy dataset
The employed data set will be formed by square images of $n\times n$ pixels, where each pixel will take values laying between 0 and 255 in grayscale. Hence, each image can be represented by a matrix $x\in GL_{n}(\mathbb{R})$. For a first stage, images with resolution $n=4$ of T and L tetrominoes [60] have been generated (see Fig. 2). Each tetromino can be placed in 6 different positions and 4 orientations inside an image of $4\times 4$ pixels, but it is possible to augment the size of the data set arbitrarily by adding a layer of noise to each image.
@@
where $x_{i,j}$ is a matrix elements representing a pixels value. A crucial step to build a label-invariant quantum model is to find the representation of this operation acting on the Hilbert space of the quantum circuit. To do so, first, the chosen data-encoding strategy will be described.
@@ III.2 Angle encoding
Frequently named as qubit encoding, angle encoding [14, 61] is one of the most simple data-encoding methods. It consists on encoding the features $x_{k}$ of the data point $x=(x_{0},\dotsi,x_{M})$ in the amplitudes of the state-vector of each qubit:
$|x\rangle=\bigotimes_{k=0}^{M}\cos(x_{k})|0\rangle+\sin(x_{k})|1\rangle.$ (10)
where $M$ is the number of features in each data point and $M=n^{2}$ for square images. This encoding can be performed by applying $R_{X}(x_{k})$ gates over each qubit (see Fig. 3), where the pixel $x_{i,j}$ is mapped to $x_{k}$ vector value with $k(i,j)=in+j$ index transformation. Due to the periodic nature of $R_{X}$ rotation gates, the classical features ought to be scaled to fit in $x_{k}\in[-\pi,\pi]$.
@@
As the rotation gates acting on different qubits commute, angle encoding can be implemented in constant circuit depth, making it one of the most fit candidates to apply on NISQ devices. This implies $n^{2}$ queries to classical memory, so the time complexity would be $\mathcal{O}(n^{2})$, where $n^{2}$ is the number of pixels (features). Regarding the number of qubits, this method requires one qubit per feature in the data-point, which can be reduced by a factor of two encoding an additional feature in the complex phase of each qubit, following the so-called dense angle encoding strategy.
This improvement in the number of required qubits $n_{q}$, which is still linear with respect to the number of features, is notably surpassed by the exponential reduction achieved with amplitude encoding [14]. Nevertheless, in this type of encoding, the algorithm proposed by Mottonen et al.[62] is usually employed for state preparation, which requires $\mathcal{O}(2^{n_{q}})$ multi-controlled gates. These gates do not commute generally, so the circuit depth of the implementation of Amplitude Encoding would be $\mathcal{O}(2^{n_{q}})=\mathcal{O}(2^{\log_{2}(n^{2})})=\mathcal{O}(n^{2})$. This depth requirements can be reduced by using ancillary qubits, and consequently loosing the advantage in the number of qubits; or encoding states with specific properties [63, 52, 53, 54].
In general, angle encoding is one of the most employed encoding strategy in QML, in part for its simplicity, both in the implementation and for theoretical constructions. For this reason, angle encoding will be employed in this work, making use of a $n_{q}=16$-qubit system, to encode images of resolution $n=4$. A way to implement the proposed algorithm for larger resolutions will be discussed in subsequent stages.
@@ III.3 Equivariant variational circuit
Figure 4: Rotation operation decomposed into a (Fig. 4(a)) transposition and (Fig. 4(b)) column permutation using SWAP operations. Example for an image of n=4.
@@
$\mathcal{T}[X_{k}]$ $=$ $\frac{1}{4}(X_{k}+U_{g}X_{k}U_{g}^{\dagger}+U_{g}^{2}X_{k}{U_{g}^{\dagger}}^{2}+U_{g}^{3}X_{k}{U_{g}^{\dagger}}^{3})$ (14)
@@
Let orbit $\varphi$ denote the set of all different indices obtained under the action of all symmetry operations for a given pair of indices $(i,j)$ (see Fig. 5(a)):
@@
where $\varphi_{k}(i,j)$ represents the component number $k$ of the orbit $\varphi(i,j)$. Consequently, the same symmetrization is obtained for rotation gates with the same rotation angle applied on qubits connected by the same orbit. This symmetrization is equivalent for the Y and Z rotation axes. The same logic can be followed for two-qubit gates, leading to this result: the same two-qubit gate must be applied over the orbits of both qubits, which must be in different orbits, to ensure the permutation condition of these gates with respect to the representation of the symmetry group. To illustrate this, as a reference to condensed matter physics, we will call unit cell to the maximal set of indices that don’t share orbits; this is, the maximal set of points that are not related by a symmetry operation (see Fig. 5(b) and Fig. 5(c)).
@@
Figure 5: 5(a)) Different orbits of an image of resolution $n=4$ denoted by colors. 5(b)) Different unit cells denoted by colors and the chosen circular connectivity of CNOT gates inside each unit cell, for an image of resolution $n=4$. 5(c)) Different unit cells denoted by colors and the chosen circular connectivity of CNOT gates between unit cells, for an image of resolution $n=4$.
Apart from the constraints concluded from the symmetrization of two-qubit gates, one has a certain freedom to choose their connectivity, i.e., on which qubit it can apply them. Indeed, the choice of the unit cell shown in Fig. 5(b) and Fig. 5(c) is not unique. A circular connectivity scheme will be selected, which entangles the qubits inside every unit cell. However, if different unit cells are not connected between them, it will produce four independent quantum circuits that would process just some few pixels of the original image. To avoid this scenario, some CNOTs can be added between unit cells and deal with a more extensive set of correlated qubits.
@@ III.4 Simulation results
An equivariant quantum circuit is trained to learn to classify tetromino images described in Sec. III.1. This circuit employs parameterized quantum gates symmetrized using the Twirling Formula, following the workflow detailed in Sec. III.3. In consequence, a label-invariant model is obtained with respect to the action of 4-fold rotations. To benchmark the power of the model and to witness whether the geometric approach does provide any advantage in comparison to conventional variational circuits, its performance is compared to a widely used architecture called Basic Entangler Layers, based on the work by Schuld, M. et al. [64]. To make the comparison the fairest possible and since the latter architecture makes use of more optimizable parameters per layer, which can induce an improved approximation capacity, general rotation gates are used in the Equivariant model, which can be decomposed into $3$ optimizable rotation gates, up to a negligible global phase [65]:
@@
It is true that these general rotation gates are more expressive than one-axis rotation gates employed in the Basic Entangler model, and this could be a possible source of a difference in performance. For this reason, a third model type is compared, which we will call Non-equivariant model. Its quantum circuit utilizes the same architecture as the Equivariant model’s but with randomly generated orbits to break the symmetrization of gates and, consequently, the equivariant condition. This way, it is ensured to capture if any improvement in the performance is due to the choice of gates or if it is, indeed, a consequence of the geometric approach.
In Tab. 1 the number of optimizable parameters per layer are shown for each model type. For the proposed equivariant architecture, the number of parameter increases linearly with the number of different orbits, which scales as $n_{\varphi}=n^{2}/4$ if $n$ is even, and $n_{\varphi}=\lceil\frac{n}{2}\rceil\lfloor\frac{n}{2}\rfloor+1$ if $n$ is odd . It is remarkable that this scaling does not stand true for the general case, since it is dependent on the gate symmetrization, determined by the chosen encoding method. For example, in the work of Chang, Su Yeon, et al. [51], and equivariant architecture was designed where it includes $6\lceil n_{q}/2\rceil+3\lceil n_{q}/4\rceil$ parameters per layer, and since amplitude encoding was employed, the scaling of the number of parameters per layer is logarithmic to the resolution of the image $n$.
Model Parameters Entangling gates
Equivariant $3n_{\varphi}$ $4(n_{\varphi}+1)$
Non-equivariant $3n_{\varphi}$ $4(n_{\varphi}+1)$
Basic Entangler $n^{2}$ $n^{2}$
@@
Table 1: The number of optimizable parameters and entangling gates (CNOT) per re-uploading layer. $n_{\varphi}$ corresponds to the number of different orbits in an image and $n$ is the resolution of the image. For an image of $n=4$, $n_{\varphi}=4$.
@@
Figure 6: Loss landscape of three model types along the training process. Ten models with randomly initialized parameters were trained for each type of architecture (Equivariant, Non-equivariant and Basic Entangler). The shaded areas are spread between the maximum and minimum costs of each model type, and the dashed lines represent the mean cost of each model type. All models apply $n_{l}=5$ re-uploading layers. These results correspond to the dataset of tetrominoes and a learning rate of $0.1$.
@@
All three models are trained in a noiseless environment using the default simulator provided by Pennylane [66], and the parameter update is performed using Adam Optimizer [67]. In Fig. 6, it can observed that the Equivariant model reaches the lowest values in the loss landscape for the same number of iterations. Even if the Basic Entangler performs a steeper descent in early steps, it reaches a plateau where the cost nearly stops decreasing. The other two models, however, keep reducing the cost until the last optimization step, although their descent tends to flatten in the last iterations. Comparing the gradients of the loss function for randomly sampled parameters, the Basic Entangler exhibits a variance two orders of magnitude smaller than that of the other two models. Similarly, the variance in loss values across different initializations is an order of magnitude smaller for the Basic Entangler model, indicating a more concentrated loss landscape compared to the other models (see Tab. 2). This aligns with the theoretical results of Ragone et al. (2023) [68]: as symmetrization reduces the dimension of the associated Dynamical Lie Algebra (DLA), the loss landscape is expected to be less concentrated for the Equivariant and Non-Equivariant models.
$\widehat{E}_{\theta}[\mathcal{L}(x,\theta)]$ $\operatorname{\widehat{Var}}_{\theta}{[\mathcal{L}(x,\theta)]}$ $\widehat{E}_{\theta}[\frac{\partial\mathcal{L}(x,\theta)}{\partial\theta}]$ $\operatorname{\widehat{Var}}_{\theta}{[\frac{\partial\mathcal{L}(x,\theta)}{\partial\theta}]}$
[…]
Table 2: Expected values and variances of the loss function and its partial derivatives for randomly sampled parameters; comparison between Equivariant, Non-equivariant and Basic Entangler models respectively. 5000 parameter sets were uniformly sampled from $[0,2\pi)$ for each type of model.
@@
Figure 7: Violin plots of F1 score metrics of trained models. Ten different models of each type are trained with the same hyperparameters ($n_{l}=5$), and the performance is evaluated over the training and test set. These results correspond to the dataset of tetrominoes and a learning rate of $0.1$
@@
Table 3: Performance metrics for each model type and per different number of data re-uploading layers. Each value in the table represents the mean performance metric of 5 different models, evaluated over the training set (left) and the test set (right). These results are for the dataset of tetrominoes.